I am unclear as to how to put this problem together:
To estimate the population of dolphins in a region, biologists capture and tag 200 dolphins, then release them. Later, they recapture 60 dolphins and find that of those captured, 4 are tagged.
a. Write an equations of ratios to solve for T, the total estimated population of dolphins.
b. What is the estimated population of dolphins in this region?
For question (a.) this is what I have:
4/60 = t/200
But when I solve it comes out to something like 13.3333333 which I know is not right. Word problems confuse me, can someone tell me what I'm doing wrong? Thank you so much!

Answer by stanbon(75887)

To estimate the population of dolphins in a region, biologists capture and tag 200 dolphins, then release them. Later, they recapture 60 dolphins and find that of those captured, 4 are tagged.
a. Write an equations of ratios to solve for T, the total estimated population of dolphins.
----
(population count)/(# caught and tagged)=(# caught including tagged)/(#tagged)
(pop ct.)/200 = (60/4)
----
b. What is the estimated population of dolphins in this region?
(pop ct.) = 200(60/4)
(pop ct.) = 3000
====================
Cheers,
Stan H.
====================